Consider palletizer at a bottling plant that has a first cost of $134144, has operating and maintenance costs of $17487 per year, and an estimated net salvage value of $55601 at the end of 30 years. Assume an interest rate of 8%. What is the future worth of this project? Enter your answer as follow: 123456.78
Answer #1

FW = P(F/P, i, n) + A(F/A, i , n) - F

       = 134144(F/P, 8%, 30) + 17487(F/A, 8% , 30) - 55601

       = 134144(10.063) + 17487(113.3) - 55601

       = 1,349,891.072 + 1,981,277.1 - 55601

       = 3,275,567.17